Initially, an object has a temperature whichhas the same numerical value in degrees Celsius and degrees Fahrenheit. Then its temperature is changed so that the numerical value
of the new temperature in degrees Celsius is 5
times as small as that in Kelvin.
Find the change in the temperature (change in T =
T[sub]2[/sub] − T[sub]1[/sub]).Answer in units of K.
1
Expert's answer
2011-02-08T08:53:03-0500
The conversion formula from Farenheight to Celsius is
[°F] = [°C] × 9/5 + 32
Denote the tempetrature in Celsius as T.
As the temperature has the same numerical value in degrees Celsius and degrees Fahrenheit
T1 = T1× 9/5 + 32, T1 = -40 [°C].

Then its temperature is changed so that the numerical value of the new temperature in degrees
Celsius is 5 times as small as that in Kelvin:
T2 = (T2+273)/5, T2 = 68.25 [°C]

Thus
T = T2 - T1 = 68.25- (-40) = 108.25 [°C].